- Any configuration change of a system needs to be approved by the Run Coordinator. (This includes removing a crate from
- Should we encounter a longer downtime during data taking and the problem does not get resolved
within 45 minutes, the Run Coordinator has to be paged.
- Cosmics -
Cosmic shifts have ended, and most systems have been turned off.
We started warming up the solenoid on Friday, Dec. 23 to conserve
helium and deal with cryo plant problems.
- SMT annealing studies are still ongoing and require special runs every one to two days.
Between the SMT bias scans DZero will be taking recording-off cosmics with the CFT/CTT
and the Muon scintillators only (i.e. no SMT, HDIs powered off).
SMT experts will take care of the necessary transitions. Please make an elog entry when interrupting datataking
for more than a few minutes, so run coordinators (who will keep an eye on the DAQ rates) are not unnecessarily alerted.
- While still using the CTT trigger, CTT examine plots should be saved to the elog and reset
before and after the SMT bias scans.
- In the second week of January 2012 we will probably end CFT/CTT operations and switch to a
Silicon-only standalone trigger.
Shift operations have ended as of noon December 23, 2011.
Global Monitoring (GM) is now performed by the shift captain.
Global Monitoring is being coordinated by
Vladimir Sirotenko and Elliot Cheu.
The GM console, d0ol104, is reserved for the GM tasks during the colliding beam operations.
Useful information, numbers and plots should be inserted into the
electronic logbook under the GM tab.
The vertex examine runs automatically (currently on d0ol77) and the Shift
Captain looks at the plots. Michiel Sanders has documented the Vertex Examine.
For reference, look at the VE plots entered in the logbook during